OpenStudy (anonymous):

I am really confused with this word problem. I don't understand what x is supposed to equal.

OpenStudy (anonymous):

Dayna, Lisa, and Lisa’s older sister, Alyssa, are trying to decide which fair they should go to next year. Fair 1 has two options, a wristband, which allows for unlimited rides and includes the cost to get into the fair, and no wristband, which has a set price to get into the fair and the cost per ride. Fair 2 has a set price to get into the fair and then a price per ticket. This year at the fair Dayna and Lisa only rode 8 rides, she is trying to decide which would be the most cost efficient option for next year. The equations showing the trend in cost for both fairs next year is displayed below. Fair #1 has two options and Fair #2 has one option. Tell which fair Dayna and her friends should choose, and why. Fair #1 wristband: y = 30x Fair #1 no wristband: y = x + 20 Fair #2: y = 1.5x + 5

OpenStudy (unklerhaukus):

x is the number of rides, and y is the total cost. to compare to last years data, consider x = 8 rides

OpenStudy (anonymous):

Well then wouldn't that mean that fair 1 with wristband would be drastically more expensive then without?

OpenStudy (unklerhaukus):

hmm, are you sure the first option isn't meant to be Fair #1 wristband: y = 30 (without the x)

OpenStudy (anonymous):

No, I typed it out exactly the way it was.

OpenStudy (unklerhaukus):

Well I think the question was written incorrectly (in your book or wherever you got the problem from)

OpenStudy (anonymous):

I think I might just answer it as y=30

OpenStudy (unklerhaukus):

Option 1: wristband, " unlimited rides and includes the cost to get into the fair", the number of rides (x) cannot increase the cost this must be of the form y = b and not y = mx+b

OpenStudy (anonymous):

Oh. That makes sense. Thanks.

OpenStudy (unklerhaukus):

Can you work it out if we assume the 'x' in option one is a typo?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

I was honestly only confused with the wording of the problem.

OpenStudy (unklerhaukus):

Fair enough, I don't like word problems much either. (especially when they have typos)

OpenStudy (anonymous):

Yeah, I was also conused whether x equaled rides or people.

OpenStudy (anonymous):

*confused

OpenStudy (unklerhaukus):

That is the only way make sense of the x in the first option, but, this interpretation doesn't really make sense with the rest of the question.

OpenStudy (anonymous):

Yeah, like you said, I will probably just answer it like y=30

OpenStudy (unklerhaukus):

That's what i would do.

OpenStudy (anonymous):

Well thanks for helping me with the problem.